OpenStudy (tatumlee):

A group of students are going on a field trip. If the group takes 3 vans and 1 car, 22 students can be transported. If the group takes 2 vans and 4 cars, 28 students can be transported. How many students can fit in each van? Can someone show me how to solve this please?

OpenStudy (phi):

let x be the number of people that fit in a van and y be the number that fit in a car then 3 vans and 1 car, 22 students means 3x + y =22 2 vans and 4 cars, 28 students means 2x +4y=28 you have two equations and two unknowns do you know how to solve this ?

OpenStudy (tatumlee):

Do I have to add the equations together some how?

OpenStudy (tatumlee):

@phi

OpenStudy (phi):

there are two ways to solve these. Substitution or elimination I usually use elimination (as that is how you solve 3 equations and 3 unknowns) 3x + y =22 2x +4y=28 the first thing to notice is that we can divide the 2nd equation by 2 (that means both sides and all terms) to get x + 2y = 14 that makes it a little simpler

OpenStudy (phi):

3x+ y= 22 x + 2y = 14 Next, I would multiply the first equation by -2 (that means both sides and all terms are multiplied by -2)

OpenStudy (tatumlee):

I am confused on where the -2 came from, and is there some sort of steps to follow in equations like these? Like first divide, then multiply then... I feel like I'm missing something

OpenStudy (tatumlee):

@phi

OpenStudy (phi):

when you have two equations like 3x+ y= 22 x + 2y = 14 and want to (for example) "get rid of the y", I notice that if the top equation had a -2y then when we add the two equations (term by term) the -2y+2y would add up to 0 (cancel out) so let's multiply the top equation by -2 (that means multiply both sides and all terms by -2) can you do that ?

OpenStudy (phi):

you should get -6x+ -2y = -44 for the 1st equation x + 2y = 14 now add term by term -5x + 0 = -30 or 5x=30 x=6

OpenStudy (tatumlee):

thank you for your help, I really don't understand but I will keep looking over what you've said and try to understand how you got where you did

OpenStudy (tatumlee):

this video is making a lot of sense to me thank you! The other videos answer some of my other questions as well.